OpenStudy (anonymous):

16x^2 + 16x = 4 Solve the quadratic equation by completing the square.

OpenStudy (anonymous):

16x2 + 16x = 4 4(4x²+4x-1)=0 x=[-4±√(16+16)]/8 x=[-1+√2]/2 x=(-1-√2)/2

OpenStudy (anonymous):

thanks you, you are a life saver

OpenStudy (anonymous):

np

OpenStudy (anonymous):

i will love it if you explain the 3rd step of the equation thanks